SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

   
Prev Next


Date: February 23, 2008
Shift Taker: Chris West
-----------
General Features:
(theta,phi)
Warm crystals in hitmap          (32,85)
Warm crystals in 0.01 GeV hitmap (17,43), (30,45), (32,46), (36,64)
Warm crystals in 0.1 GeV hitmap  (30,45), (32,46), (36,64)
Warm crystals in 1 GeV hitmap      

Cool crystal  in hitmap          
Cool crystal  in 0.01 GeV hitmap (29,84) 
Cool crystal  in 1 GeV hitmap    (3,43)

Masked ADB 3-1-4.
Runs with >1e-3 OOS/OOO damage in 0-5: 79836, 79840, 79841, 79843, 79844, 79853
79827-79828: OK
79829: Strange hits in  in 0.1 GeV hitmap theta> 29, phi~=45-68, background spikes
79830: OK
79831: Hints of new type of strange hits.
79832-79833: OK
79834: >1e-3 OOS in 4-10
79835: OK
79836: >1e-3 OOO/OOS in 4-10
79837-79846: OK.
79847: Blobs at t~6650ns E~1-13 GeV (backgrounds)
79848: OK
79849: >1e-3 OOS in 1-4
79850: Timing blobs from background
79851: OK
79852: High HER backgrounds
79853: OK
79860-79862: OK


Crystal Spectra:
"Badness 1" at theta=37-40, phi=43

Lightpulser:
16 channels dead, 0 channels out of time. Updated on: 23-February 2008 21:24 PST



 

  Links

  • Shift Report Index
  • Back to the Shifter instructions